OpenStudy (timaashorty):

simplify the expression. -6+i/-5+i

OpenStudy (timaashorty):

THESE ARE THE CHOICE ANSWERS: \[A.\frac{ .31+i }{ 26}\] \[B.\frac{ 29+i }{ 26}\] \[C.\frac{ .31+i }{ 24}\] \[D.\frac{ .31+11i }{26}\]

OpenStudy (anonymous):

-6 + 0.8 i

OpenStudy (anonymous):

multiply with i+5 on both numerator and denominator

OpenStudy (anonymous):

@timaashorty are you listening

OpenStudy (timaashorty):

@thisSucks I'm sorry I wasn't on the computer. Anways I believe the answer is A. Is that correct ?

OpenStudy (anonymous):

NO

OpenStudy (timaashorty):

How is it not correct ? can you help me figure the right answer ?

OpenStudy (cggurumanjunath):